# Learning to Visually Connect Actions and their Effects

### v1

@@ Abstract
In this work, we introduce the novel concept of visually Connecting Actions and Their Effects (CATE) in video understanding. CATE can have applications in areas like task planning and learning from demonstration. We propose different CATE-based task formulations, such as action selection and action specification, where video understanding models connect actions and effects at semantic and fine-grained levels. We observe that different formulations produce representations capturing intuitive action properties. We also design various baseline models for action selection and action specification. Despite the intuitive nature of the task, we observe that models struggle, and humans outperform them by a large margin. The study aims to establish a foundation for future efforts, showcasing the flexibility and versatility of connecting actions and effects in video understanding, with the hope of inspiring advanced formulations and models.
@@ 1 Introduction
[…]
[loop,autoplay,poster=1,width=0.9]3Figs/fig_concept_frs/fr_0112
Figure 1: Concept. Please view in AdobeReader to play the embedded videos for better explanation. We have shown two exemplary questions, from our task, alongwith candidate action options. Correct answers are (b) (not (c)) & (a) in questions 1 & 2, resp.
Actions can transform systems, scenes, and environments. Humans can connect their actions with the corresponding effects, i.e., given an initial state [^1] In general, the state of the system before the action is applied is termed as initial state, and the state of the system after the action is applied is termed the final state., when some action is applied, we can imagine the outcome or the resultant state of the system—e.g., consider a scenario where a person is holding an object; if they release their grip, we intuitively know that the object will inevitably fall to the ground. This understanding of causal mechanisms is crucial as it allows humans to plan and execute goal-directed actions. Humans can anticipate the consequences of their actions, make informed decisions, and adjust their behavior. The integration of visual information with motor actions allows individuals to build a coherent and adaptive representation of the world, enabling them to navigate their environment effectively. The presence of such capabilities in video understanding systems and autonomous robots is essential, with applications spanning areas such as task planning and learning from demonstration. This not only makes them versatile but also enables real-time adjustments to their behavior. Despite being recognized as a fundamental aspect of human cognition, the capability to link actions and their physical effects remains largely unexplored in the existing body of literature on video understanding. Towards that end, in this work, we introduce a novel concept of visually connecting actions and their effects.
@@
Connecting actions and their effects can be approached from various perspectives, leading to different task formulations. For instance, actions and effects could be linked at a semantic level. Following this perspective, we introduce the task of Action Selection, which requires video understanding models to visually connect actions and their effects. In this task, a video understanding model is presented with a pair of initial and final states of a scene, and it must select the correct action from a set of options that transitions the system from the initial state to the final state, as illustrated in Figure 1.
@@
Now, lets take a look at connecting actions and effects at a finer granularity. To achieve this, we introduce the task of Action Specification, where a video understanding model is tasked with inferring how closely or directly related is an effect to an action. So, while Action Selection requires connecting actions and effects at a semantic level with coarser granularity, Action Specification involves discerning at a finer resolution.
@@
Interestingly, we observe that different formulations of connecting actions and their effects learn representations that capture intuitive properties of actions. For example, the Action Selection model learns to track without any explicit supervision (see Figure 5); while Action Specification representations learn to encode the pose of humans without any supervision (see Figure 8). We also observe that connecting actions and effects can serve as an effective self-supervised pretext task to learn video representations from unlabeled videos. Furthermore, we note that the ability to generate effect can benefit other tasks such as action anticipation.
While connecting actions and their effects may seem an easy task for humans, we found it to be quite challenging for state-of-the-art video understanding models. To this end, we propose several baseline frameworks which improve over the current state-of-the-art video understanding models, but we find that there is still a large performance gap between humans and machines on this task.
We aim to lay a solid foundation for future efforts through devising multiple task formulations and models. Multiple formulations show that connecting actions and their effects is a flexible and versatile task with many possible applications. Developing and evaluating diverse set of baseline models/solutions helps ensure comprehensive coverage of the problem space and provides a more robust evaluation framework. We hope that these will inspire future efforts to develop more advanced formulations and models with further applications and better performance.
The paper is organized as follows. We start by reviewing related work in Section 2, followed by task formulations in Sections 345. For better organization, Models and Dataset are discussed along with their corresponding task. Task and models are evaluated and validated in Section 6; finally, concluding the paper in Section 7.
@@ 2 Related Work
Video understanding is a mature branch of computer vision addressing action analysis problems such action recognition [35, 3, 10, 39, 25], action segmentation/localization [19, 5, 6, 33], and action quality assessment [29, 31, 47, 1]. While tremendous progress has been made in these directions, connecting actions and their effects largely remains to be addressed. Towards that end, in this work, we study linking actions and their effects and study consequences of the same. We make use of the techniques and insights gleaned from existing video understanding tasks towards the problem of connecting actions and effects.
Viewing actions as transformations.
@@
Actions can be viewed as transformations or agents of transformations occurring in the scene or an environment [42, 36, 27]. Based on this, Wang et al. [42], proposed to model actions as fixed class-specific transformation matrices, which when applied to a precondition would produce the effect. Soucek et al. [36] propose to learn action-specific models to localize states and actions. Similar to actions, Nagarajan and Grauman [27] view object properties as operators on objects. However, these work either focus on learning a rigid set of class-specific transformation matrices [42], or learning only action-specific models [36] or limited to static images [27, 36] and do not focus on studying action dynamics and their effects. Our work studies linking action dynamics and effects in much more depth and breadth. Our task also demonstrates wider applicability. We further take inspiration of such prior work and develop their dynamic counterparts which can generate transformations on the fly from actions and that way help in connecting actions and their effects.
Synthetic datasets.
Obtaining the groundtruth for object properties and positions can be made easier and efficient with the help of synthetically generated datasets [18, 9]. As such, synthetic datasets form a lucrative option for studying state-changes as done in [15]. However, sophisticated approaches developed towards such synthetic datasets are not generally not applicable to real-world scenarios where the models do not have access to high level of control over the attributes of states and actions. Our work conducts all the studies and validating experiments on real-world datasets. Furthermore, our novel cross-sample setting can be seen as a real-world equivalent to View transformation [15] or even more challenging as many other factors are varied. Another difference between our work and those using synthetic data is that they do not contain actions involving humans, while ours has been experimented specifically with human actions in the real-world.
Modeling state-changes.
@@
A series of work [28, 17, 20] address the problem of computing state changes from the initial and final states of the environment. Computing state-changes is an important part of our proposed problem. However, unlike our work, these work do not link the changes in state with the actions driving those changes.
Assorted work on visual reasoning.
@@
Liang et al. [22] study an interesting problem of coming up with likely explanations for partial observations. Parallels could be drawn between initial and final states and partial observations. However, their work is concerned with forming semantically higher-level explanations, whereas ours is more focused on foundational level connection—linking action mechanics or kinematics with their effects. Other interesting reasoning problems have been introduced in recent years [43, 13], but they are orthogonal to ours.
3 Task Formulation-I: Action Selection
Task formulations are crucial as each one has a unique objective and enables control over the model’s focus to learn representations that capture slightly different types of visual features. We start by discussing the first task formulation—Action Selection.
Objective.
Given the initial and desired final states, this task formulation requires video understanding models to select the correct action instance that changes the scene from its initial to its final state. Taking inspiration from visual question-answering literature, we cast the problem as a multiple-choice question answering; and provide four action-video choices for answers—see Figure 1. From a total of four choices/options, only one is correct. The task is to select this correct action-video. Action Selection task can enable Autonomous Agent/Robot Action Selection from human demonstrations from existing large-scale video datasets—discussed in subsection 3.1.
Constructing Cross-Sample question and answer sets.
@@ 3 Task Formulation-I: Action Selection
To form a question, we divide a video instance of a human performing an action into the initial state of the scene (starting few frames), and the resultant final state (last few frames). The remaining middle frames after keeping a certain margin from the initial and final state frames become the correct action option. However, the correct answer option can be drawn in two ways: 1) same-sample setting: from the same video clip as the states as explained earlier; or 2) cross-sample setting: from an altogether different instance, but the action category still remains the same as the video from which states are drawn. Cross-sample setting can be viewed as a very strong data augmentation because essentially all factors such as background, colors, viewing angle, actors, and scene setup may be different from state-video, while semantic action class remains the same. This setup enables the isolation of action semantics from irrelevant features of the scene and demands to focus on the relevant parts of the video, thereby helping connect actions and their effects effectively.
3.1 Dataset
@@
We introduce a novel dataset to support our proposed Action Selection task. Each sample in our dataset consists of the “question”, and the corresponding a single correct answer and three incorrect answers.
Dataset sources
@@
The first challenge in creating our dataset was choosing appropriate large-scale video datasets containing numerous action classes involving significant state changes. For this we chose: 1) Something-Something version 2 (SSv2) dataset [10]; and 2) COIN dataset [38]. SSv2 encompasses a diverse range of actions — physical movements, object handling, transformations, and interactions. It contains actions related to the manipulation of physical objects, such as lifting, pushing, pouring, stacking, and more, as well as actions that involve altering the position and shape of objects. COIN dataset on the other hand primarily centers on the execution of specific tasks and procedures involved in assembling, installing, building, or manipulating various objects and materials. It often comprises step-by-step instructions for tasks like assembly, installation, cooking, and maintenance, with an emphasis on precise and ordered procedures, including actions like “installing,” “cutting,” “putting,” “removing,” and “filling”. Overall, using SSv2 and COIN helps us build a very diverse dataset.
Selecting action classes.
Not all action classes from SSv2 and COIN involve significant state changes, and consequently, might be less suitable for CATE Action Selection task. For example, action class like “approaching something with camera” is less likely to involve intended state changes; and as such, should be discarded. To mitigate, we manually filter usable action classes from these base datasets. To determine if an action class is usable, we first take into consideration the name of action class. We further inspect randomly sampled video from the action classes to verify if the action class contains significant intended state changes. Based on these two evaluations, we decide if the action class is usable. Overall, a combined 204 action classes are found to be usable. We further ensure that there is no overlap between action classes to avoid confusing the models.
Temporal annotations for states and actions.
@@
To effectively capture states and actions, we utilize the first and last 10% of all video frames as the initial and final states, respectively. The middle 80% represents the driving action, allowing flexibility for varying video lengths using percentages rather than fixed frame counts. Given potential noise in crowd-sourced datasets like SSv2, where clips may include irrelevant or shaky footage, we address this by detecting objects and hands. We retain video footage only after hands or relevant objects are detected and stop retention when they are not detected. This approach enhances dataset quality, as illustrated in supplementary.
Constructing quadruplets.
@@
We follow the cross-sample setting as mentioned in section 3 to construct quadruplets of question, correct answer, and incorrect answer options.
Causal-Confusion Hard Counterfactuals.
@@
Incorrect action options can be created by simply choosing drawing them from altogether a different action category than the category of the correct action option. While this strategy suffices the requirement, we hypothesize that harder to distinguish incorrect options can be generated. Towards this end, we introduce the Causal-Confusion hard counterfactual/incorrect action option generation strategy. For this we take the correct action option and apply one of the following transformation: 1) temporal-reversal—temporally reverse the action clip; 2) static—repeat the first action frame, which causally means that no real action is carried out; 3) flipping the video horizontally—actions moving something left to right will have completely opposite meaning. Note that with this strategy, samples can be generated that look identical to the correct action class, but differ in terms of causal factors. As such, this strategy encourages the video understanding model to focus on causal factors, while learning to be invariant to irrelevant factors such as background and appearance.
3.2 Baseline models
Figure 2: Models. Please zoom in for a better view.
As the proposed task is new, and there are no established methods or benchmarks available for tackling this specific task, we design and explore a wide variety of baseline methods. For simplicity, only the best-performing methods are presented here, remaining methods are included in supplementary material.
Preliminaries.
  • Action & State Representations. Actions are represented as video clips, while states by very short clips. Unless specified otherwise, we use a pre-trained backbone model from a large-scale action recognition dataset to extract features for actions and states. Subsequently, various modules mentioned in the following are trained on these features.
  • Similarity vs Counterfactual reasoning. All of our methods involve feature matching in some form during training. Matching could be done just based on the similarity with the correct output or done via counterfactual reasoning by using similarity along with contrasting with incorrect outputs. In our experiments, we found counterfactual reasoning to work significantly better than similarity-based reasoning. Thus, unless mentioned otherwise, assume that counterfactual reasoning is employed
Naive matching.
As a simple baseline, in naive matching the average of the initial and final state features are matched with the action features using cosine similarity. The action answer with the highest similarity score is selected as the correct answer. This model involves no training, pretrained action recognition model is used as the state and action feature extractor.
Treating actions as transformations.
Actions can be viewed as transformations [42] which when applied to the initial state yields the final state. While the original approach uses fixed transformations [42], our adaptation facilitates conditional transformations computed on the fly. We develop a model in which a transformation vector is computed from an action clip ($I$) using a dense network. This transformation vector or matrix ($A$) when multiplied with the initial state vector yields the final state vector $F$ (Figure 2(a)). During training, we use contrastive learning [4], where the cosine similarities between the predicted and wrong action options are minimized while the similarity between predicted and correct action options is maximized. During inference, the action that produces the highest similarity with the final state feature is selected as the correct answer. We term these models as Actions as transformations (AT).
Modeling rich interactions between initial state & action (MoRISA).
Here a bilinear model [23] is leveraged to capture rich interactions between the initial state and the action to generate the final state in feature space. Essentially, initial state ($I$) and action representations ($A$) are fed into a bilinear model, which outputs a final state representation ($F$) (Figure 2(b)). The parameters of the model are trained using contrastive learning as before and the inference is done similarly to AT model above.
Learning linear embedding State-Action space.
In this framework, actions and initial states undergo transformation into a linear embedding state-action representation space through a Transformer encoder. The hypothesis suggests that within this linear, disentangled space, adding an action vector to the initial state representation enables a transition to the final state in the representation space, as depicted in Figure 2(d). The Transformer encoder processes a sequence of initial state and action representation vectors, and the resulting output is decoded by a linear decoder to produce the final state feature vector. During training, all model parameters are optimized end-to-end with the objective of aligning the resultant final state representation with the ground truth final state representation. Optimization and inference process remains same as AT model. We call this model AEXformer.
Explicitly Disentangling State-Transformation.
To explicitly isolate a $n$-dimensional “transformation “code”, we pass the initial and final states (each represented by $m$-dimensional features) through an encoder, which outputs a $(m+n)$-dimensional vector (Figure 2(c)). We enforce the condition that the first $m$ elements of this vector represent the initial state and the remaining $n$ elements represent the transformation “code”. Now, using another encoder, we distill a $n$-dimensional transformation code from an action option. We swap the transformation code obtained using states with that obtained from the action option. A decoder then reconstructs the final state from the extracted initial state and swaps the transformation code. All encoders and decoders are trained by enforcing that the extracted initial state be similar to the ground truth initial state and the reconstructed final state be similar to the ground truth final state in case the transformation code was coming from the correct action option (minimize the similarity if the transformation code was from incorrect action option). This approach is inspired from [31]. However, while their approach disentangles human pose and appearance, our approach extends it to disentangle state-transformation for linking actions and their effects.
Analogical Reasoning.

### v3

@@ Abstract
We introduce the novel concept of visually Connecting Actions and Their Effects (CATE) in video understanding. CATE can have applications in areas like task planning and learning from demonstration. We identify and explore two different aspects of the concept of CATE: Action Selection (AS) and Effect-Affinity Assessment (EAA), where video understanding models connect actions and effects at semantic and fine-grained levels, respectively. We design various baseline models for AS and EAA. Despite the intuitive nature of the task, we observe that models struggle, and humans outperform them by a large margin. Our experiments show that in solving AS & EAA, models learn intuitive properties like object tracking and pose encoding without explicit supervision. We demonstrate that CATE can be an effective self-supervised task for learning video representations from unlabeled videos. The study aims to showcase the fundamental nature and versatility of CATE, with the hope of inspiring advanced formulations and models. Dataset and Code will be publicly released.
@@ 1 Introduction
Actions can transform systems, scenes, and environments. Humans possess the remarkable ability to connect actions with their corresponding effects (CATE) and vice versa based on visual information. Given an initial visual state [^1] In general, the state of the system before the action is applied is termed as initial state, and the state of the system after the action is applied is termed the final or the desired state., we can select an appropriate action and execute it to achieve the desired state/outcome, e.g., as in Figure 1. Conversely, given an initial state, when an action is applied, we can imagine the outcome or the resultant state of the system—for instance, consider a scenario where a person is holding an object; if they release their grip, we intuitively know that the object will inevitably fall to the ground.
This bidirectional understanding of causal mechanisms is crucial as it allows humans to plan and execute goal-directed actions, anticipate the consequences, make informed decisions, and adjust their behavior. Integrating visual information with motor actions helps individuals build a coherent and adaptive representation of the world, enabling effective navigation of their environment. Incorporating such capabilities into video understanding systems and autonomous robots is essential, with applications in task planning and learning from demonstration. CATE can enable practical applications like: 1) Robots can learn to perform complex tasks by understanding the relationships between their actions & the resulting effects in the environment; 2) Sports AI coach & rehab by analyzing athletes’ current movements & ideal movements, AI system can suggest personalized training plans; 3) Education: Students can interact with CATE-powered simulations to understand how different actions lead to specific outcomes in physics, biomechanics, engineering.
Figure 1: Ability to select the action to carry out to achieve the desired results is so effectively and widely used by humans, that it has become second nature to us—we use it without even realizing we are using. But can the video understanding models do the same thing? Correct answer is (c), not (b).
Despite its fundamental role in human cognition, linking actions to their physical effects is largely unexplored in video understanding. To address this, we introduce the concept of visually connecting actions and effects. Our work contributes to this understanding by investigating two key aspects of CATE: 1) Action Selection and 2) Effect-Affinity Assessment, offering distinct perspectives on connecting actions to effects in visual scenes.
@@
Connecting actions and their effects can be approached from various perspectives, leading to different task formulations. For instance, actions and effects could be linked at a semantic level. Following this perspective, we introduce the task of Action Selection, where video understanding models are presented with a pair of initial and final visual states of a scene, and they must select the correct action from a set of options that transitions the system from the initial state to the final or desired state (ref. Figure 1). While Action Selection may seem an easy and intuitive task for humans, we found it to be quite challenging for various state-of-the-art video understanding models. To address this, we propose several baseline frameworks that improve upon the current state-of-the-art models, but there remains a significant performance gap between humans and machines in this task.
@@
Now, let us take a look at connecting actions and effects at a finer granularity. To achieve this, we introduce the task of Effect-Affinity Assessment, where a video understanding model is tasked with inferring how closely or directly related is an effect to an action (Figure 1, Figure 4). So, while Action Selection requires connecting actions and effects at a semantic level with coarser granularity, Effect-Affinity Assessment involves discerning at a finer resolution.
@@
Interestingly, we observe that different formulations of CATE learn representations that capture intuitive properties of actions. For example, the Action Selection model learns to track objects without any explicit supervision (see Figure 5); while Effect-Affinity Assessment representations learn to encode fine-grained details like the pose of humans without explicit supervision (see Figure 7). We also observe that connecting actions and effects can serve as an effective self-supervised pretext task to learn generic video representations from unlabeled videos.
We aim to lay a foundation for future work by exploring various aspects of CATE, devising task formulations, & developing models. Our study underscores CATE’s essential role in learning, highlighting its flexibility across contexts & versatility across domains. This paves the way for advanced models with broader applications & improved performance.
The paper is organized as follows: §2 reviews related work. §3 discusses Action Selection task, dataset & baseline model. §4 shows that useful representations emerge from Action Selection. §5 covers Effect-Affinity Assessment & its model. §6 benchmarks baselines on CATE tasks & examines learnt representations. §7 concludes the paper.
@@ 2 Related Work
Video understanding tasks. Video understanding is a mature branch of computer vision addressing action analysis problems such action recognition [42, 3, 12, 47, 28], action segmentation/localization [21, 5, 8, 39], and action quality assessment [35, 34, 55, 1]. Tremendous progress has been made in these directions, but connecting actions & their effects largely remains to be addressed. Towards that end, in this work, we study linking actions & their effects and study consequences of the same. We make use of the techniques & insights gleaned from existing video understanding tasks towards the problem of connecting actions & effects.
@@
Viewing actions as transformations. Actions can be viewed as transformations or agents of transformations occurring in the scene or an environment [50, 43, 31]. Based on this, Wang et al. [50], proposed to model actions as fixed class-specific transformation matrices, which when applied to a precondition would produce the effect. Soucek et al. [43] proposed to learn action-specific models to localize states and actions. Similar to actions, Nagarajan and Grauman [31] view object properties as operators on objects. However, these works either focus on learning a rigid set of class-specific transformation matrices [50], or learning only action-specific models [43] or limited to static images [31, 43] and do not focus on studying action dynamics and their effects. Our work studies linking action dynamics and effects in much more depth and breadth. Our task also demonstrates wider applicability. We further take inspiration from prior work & develop their dynamic counterparts which can generate transformations on the fly from actions & that way help in connecting actions & their effects.
Synthetic datasets. Synthetic datasets simplify obtaining ground-truth for object properties and positions [20, 11] and are useful for studying state changes [17]. However, these methods are less applicable to real-world scenarios, which lack control over state and action attributes. Our work uses real-world datasets and introduces a cross-sample setting, akin to View transformation [17] but more challenging due to varied factors. Unlike synthetic datasets, our experiments involve human actions in real-world contexts.
@@
Modeling state-changes. Computing state changes from initial & final states is addressed in works like [33, 19, 22]. While this is crucial to our problem, these works do not link state changes to the actions causing them, unlike our work.
@@
Assorted work on visual reasoning. Liang et al. [25] address generating likely explanations for partial observations, which is related to our work on linking initial and final states. However, they focus on higher-level explanations, while we focus on foundational connections between action mechanics and effects. Other reasoning problems introduced recently [51, 16] are orthogonal to ours.
3 Problem of Action Selection
Given the initial & desired final states, the problem of Action Selection involves video understanding models to select the correct action instance that can achieve the desired state. In practical scenarios, we have initial state, but the final desired state may not be directly observable but exists as a mental representation (or other forms such as a sketch). However, to create a simplified/proof-of-concept version of the task, we assume access to the initial & final states in the form of images. This approach helps isolate & focus on the core challenge of understanding causal relationships between actions & outcomes without the additional complexity of environmental modeling.
Task Design. We cast the problem of Action Selection as a multiple-choice question answering; and provide four action-video choices for answers (similar to  Figure 1—we have shown only three options there for simplicity). From a total of four choices/options, only one is correct. The task is to select the correct action-video. The multiple-choice format for the Action Selection task is not only practical and reflective of both human cognitive processes and real-world autonomous agent operational scenarios, but it also aids in simplifying and standardizing the learning and evaluation processes. In practical scenarios more than four action choices/options are expected. However, as we will see in experiments, the performance of models was already quite low even with just 4 options; with more number of options the performance is only expected to go down from there. Thus, we did not consider more options in this pilot study.
Looking ahead, the Action Selection task has the potential to enable autonomous agents and robots to learn action selection from human demonstrations found in the existing large-scale video datasets, as discussed in subsection 3.1.
@@ 3 Problem of Action Selection
Constructing Cross-Sample question-answer sets. To form a question, we divide a video instance of a human performing an action into the initial state of the scene (starting few frames), & the resultant final state (last few frames). The remaining middle frames after keeping a certain margin from the initial & final state frames become the correct action option. However, the correct answer option can be drawn in two ways: 1) same-sample setting: from the same video clip as the states as explained earlier; or 2) cross-sample setting: from an altogether different instance, but the action category remains the same as the video from which states are drawn. Cross-sample setting can be viewed as a very strong data augmentation because essentially all factors such as background, colors, viewing angle, actors, & scene setup may be different from state-video, while semantic action class remains the same. This setup enables the isolation of action semantics from irrelevant features of the scene & demands to focus on the relevant parts of the video, thereby helping connect actions & their effects effectively.
3.1 Cross-Sample Action Selection Dataset
@@
We introduce a novel dataset to support our proposed Action Selection task. Each sample in our dataset consists of the “question” (pair of initial and desired states), and the corresponding a 1 correct answer and 3 incorrect answers.
@@
Dataset sources. To ensure a diverse dataset, we chose large-scale video datasets with numerous action classes involving significant state changes: 1) SSv2 dataset [12] encompasses a diverse range of actions, including physical movements, object handling, transformations, and interactions. It includes actions such as lifting, pushing, pouring, stacking, and more, involving the manipulation and alteration of objects. 2) COIN dataset [45] focuses on specific tasks and procedures related to assembling, installing, building, or manipulating various objects and materials. It includes step-by-step (we consider a step as an action) instructions for tasks like assembly, installation, cooking, and maintenance, emphasizing precise and ordered procedures.
Selecting action classes. Not all action classes from SSv2 and COIN involve significant state changes, making them less suitable for the Action Selection task. For instance, classes like “approaching something with camera” were discarded. We manually filtered usable action classes by evaluating their names and inspecting randomly sampled videos to verify significant state changes. This process yielded a total of 204 usable action classes. No overlap of classes was ensured manually to avoid model confusion. Further details on dataset are provided in Appendix.
@@
Temporal annotations for states and actions. To capture states and actions effectively, we use the first & last 10% of all video frames as the initial and final states, respectively, with the middle 80% representing the driving action. Using percentages rather than fixed frame counts, this method accommodates varying video lengths. To address potential noise in crowd-sourced datasets like SSv2, where clips may include irrelevant or shaky footage, we detect objects and hands, retaining footage only when they are detected. This enhances dataset quality, as shown in Appendix.
@@
Constructing quadruplets. We follow the cross-sample setting as mentioned in section 3 to construct quadruplets of question, correct answer, & incorrect answer options. In total, we prepare 2 million question-answer sets (1M from each SSv2, COIN datasets). Number of Train/Val/Test questions: 1.4M, 0.3M, 0.3M.
@@
Causal-Confusion Hard Counterfactuals. Incorrect action options can be created by drawing them from different action categories. However, we introduce a harder-to-distinguish strategy called Causal-Confusion hard counterfactuals. This involves applying one of the following transformations to the correct action option: 1) Temporal-reversal: Temporally reverse the action clip (e.g., option (b) in Figure 1). 2) Static: Repeat the first action frame, meaning no real action occurs. 3) Horizontal flipping: Actions moving left to right will have opposite meanings. These transformations generate samples that look identical to the correct action class but differ in causal factors. We manually shortlisted appropriate augmentations for each action, ensuring that unsuitable augmentations are not applied to any action. The augmentation list will be released along with the dataset to aid future work. This strategy encourages the video understanding model to focus on causal factors while learning to be invariant to irrelevant factors such as background and appearance.
Figure 2: Cross-sample Analogical reasoning model. Here, we have shown only one incorrect action; in practice, we use all the incorrect options for counterfactual reasoning.
3.2 Cross-Sample Analogical Reasoning Model for Action Selection
Inspired by [32, 15], we develop a human cognition-inspired baseline model for Action Selection—Analogical Reasoning model, detailed below. Model is shown in Figure 2. Model overview. The goal of the Analogical Reasoning model is to select the correct action that can help achieve the desired state starting from the initial state. During training, the model learns to map state changes & actions into a joint space using contrastive learning [4]. This involves maximizing the (cosine-)similarity between the representations of state changes & correct actions while minimizing the similarity between the representations of state changes & incorrect actions. During inference, the alignment between state change & action representations is measured in this joint space. The action with the best alignment (cosine similarity) to the state change feature vector is selected as the correct action from the available action options to achieve the desired state from the initial state.